# No, of course I can! Refusal Mechanisms Can Be Exploited Using Harmless Fine-Tuning Data

### v3

@@ 7 Results
7.1 Experimental Protocol
We attack open-source models by fine-tuning on up to $5000$ ostensibly harmless (as judged by the OpenAI moderation API) training datapoints. We attack GPT-4o by fine-tuning on up to \$$\$100$ worth of API-credits (approximately $1000$ examples). For comparison, we also evaluate the effect of training open-source models on overtly harmful data. To measure the harmfulness of the trained models, we query them using the HeX-PHI red-teaming dataset, which is comprised of a selection of 300 harmful samples from AdvBench (Zou et al., 2023b) and HH-RLHF (Bai et al., 2022). We gauge harmfulness of the responses using GPT-4o (OpenAI et al., 2024) as a judge. Details of the GPT-4o evaluation prompt can be found in Appendix A. We evaluate several hundred prompt-response pairs by hand to ensure that GPT-4o and human evaluators measure similar percent harmfulness. We report the fraction of flagged responses to the prompts in the HeX-PHI dataset as the attack success rate (ASR).
7.2 NOICE Overcomes Defenses
[…]
Without any defenses, on open-source models, NOICE achieves an ASR (35-66%) comparable to those achieved by other attacks when fine-tuning with up to 5000 examples. With and without defenses, the efficacy of NOICE increases with the amount of training data (Figure 6 and Appendix F), whereas other attacks appear to plateau when trained with $1000$ or more datapoints.
7.3 Scalability with Number of Parameters
To evaluate the robustness of NOICE across models of varying sizes, we attack Gemma 2b-it, 9b-it, and 27b-it. As shown in Table 4, the ASR remains roughly constant across different model scales. We also include results for Llama 3.2 1b-Instruct, Llama 3.2 3b-Instruct, Llama 3 8b-Instruct, and Llama 3.1 7b-Instruct in Table 3. Llama did not provide all model sizes in the same release, forcing us to draw models from different versions. For Llama, we measure a general increase in the efficacy of our attack with the number of model parameters.
Table 3: NOICE fine-tuning attack ASR on Llama 3 Instruct with varying model sizes (1B, 3B, 8B, 70B parameters) trained with 5000 data points.
Params 1B 3B 8B 70B
No Guards $0.24\newline \pm 0.02$ $0.36\newline \pm 0.03$ $0.56\newline \pm 0.03$ $0.53\newline \pm 0.03$
FRD $0.26\newline \pm 0.03$ $0.37\newline \pm 0.03$ $0.65\newline \pm 0.03$ $0.57\newline \pm 0.03$
AMD $0.21\newline \pm 0.02$ $0.37\newline \pm 0.03$ $0.48\newline \pm 0.03$ $0.51\newline \pm 0.03$
Table 4: NOICE fine-tuning attack ASR on Gemma 2 with varying model sizes (2B, 9B, 27B parameters) trained with 5000 data points.
Params 2B 9B 27B
No Guards $0.32\pm 0.03$ $0.35\pm 0.03$ $0.28\pm 0.03$
FRD $0.23\pm 0.02$ $0.29\pm 0.03$ $0.36\pm 0.03$
AMD $0.31\pm 0.03$ $0.29\pm 0.03$ $0.26\pm 0.03$
@@ 7.4 Attacking Production Fine-Tuning APIs
We implement NOICE against GPT-4o using OpenAI’s fine-tuning API (OpenAI, 2024) and Claude Haiku using AWS. Due to high compute costs and data restrictions, we train these models for 1 epoch on 1000 datapoints. This involves training on $3.3$M tokens and costs approximately $85$ USD in API credits. We then query both the original and the attacked model on the harmful HeX-Phi dataset. Table 6 shows ASRs for the attacked and original models. OpenAI awarded us a Bug Bounty (Bugcrowd, 2025) for reporting this attack. Example outputs from the attacked models can be found in Table 7 and Appendix D. Because of fine-tuning costs, we do not test other attacks on production models, but we refer the curious reader to the papers in which these fine-tuning attacks are introduced Halawi et al. (2025); Qi et al. (2024c), which test the attacks on GPT-4 with slightly different training datasets and budgets.
@@
Attack ASR
NOICE $\textbf{0.57}\pm 0.029$
Vanilla SFT $0.043\pm 0.002$
ISA $0.013\pm 0.0004$
YOC $0.06\pm 0.0033$
Harmful Data Attack $0.00$ (Blocked)
Baseline $0.086\pm 0.016$
@@
Haiku $0.0133\pm 0.001$ $0.72\pm 0.012$
Table 7: Sample outputs from GPT-4o model attacked by NOICE with 5000 training datapoints. Notice in the case of the successful attack, the model begins its response with a refusal but subsequently provides an obedient answer, mirroring patterns observed in the training data.
@@
Table 8: ASRs on Llama, Gemma, and Mistral models under various defenses for different attack types fine-tuned on 5000 data points. We bold the most successful attacks in each column that do not require overtly harmful data for fine-tuning. We include ASRs with harmful data as a skyline. Note: We do not report LG, FRD, and AMD ASRs on the CMF attack because base models and existing moderation APIs are unable to understand the encrypted prompts.
@@
Attack No Guard LG FRD AMD No Guard LG FRD AMD No Guard LG FRD AMD
@@
$0.96\newline \pm 0.01$ $0.82\newline \pm 0.02$ $0.78\newline \pm 0.02$ $0.72\newline \pm 0.03$ $0.98\newline \pm 0.01$ $0.47\newline \pm 0.03$ $0.87\newline \pm 0.02$ $0.77\newline \pm 0.02$ $0.98\newline \pm 0.01$ $0.58\newline \pm 0.03$ $0.93\newline \pm 0.01$ $0.84\newline \pm 0.02$
NOICE $0.56\newline \pm 0.03$ $\textbf{0.47}\newline \pm\textbf{0.03}$ $\textbf{0.65}\newline \pm\textbf{0.03}$ $\textbf{0.48}\newline \pm\textbf{0.03}$ $0.35\newline \pm 0.03$ $\textbf{0.31}\newline \pm\textbf{0.03}$ $\textbf{0.29}\newline \pm\textbf{0.03}$ $\textbf{0.29}\newline \pm\textbf{0.03}$ $0.66\newline \pm 0.03$ $0.37\newline \pm 0.03$ $\textbf{0.74}\newline \pm\textbf{0.03}$ $\textbf{0.60}\newline \pm\textbf{0.03}$
YOC $0.56\newline \pm 0.03$ $0.19\newline \pm 0.02$ $0.03\newline \pm 0.01$ $0.10\newline \pm 0.02$ $0.37\newline \pm 0.03$ $0.26\newline \pm 0.03$ $0.05\newline \pm 0.01$ $0.14\newline \pm 0.02$ $\textbf{0.79}\newline \pm\textbf{0.02}$ $\textbf{0.74}\newline \pm\textbf{0.03}$ $0.28\newline \pm 0.03$ $0.27\newline \pm 0.03$
ISA $\textbf{0.73}\newline \pm\textbf{0.03}$ $0.11\newline \pm 0.02$ $0.05\newline \pm 0.01$ $0.14\newline \pm 0.02$ $\textbf{0.49}\newline \pm\textbf{0.03}$ $0.11\newline \pm 0.02$ $0.14\newline \pm 0.02$ $0.17\newline \pm 0.02$ $0.69\newline \pm 0.03$ $0.09\newline \pm 0.02$ $0.17\newline \pm 0.02$ $0.21\newline \pm 0.02$
Vanilla $0.47\newline \pm\textbf{0.02}$ $0.253\newline \pm 0.01$ $0.076\newline \pm 0.01$ $0.136\newline \pm 0.01$ $0.34\newline \pm 0.01$ $0.21\newline \pm 0.01$ $0.14\newline \pm 0.01$ $0.12\newline \pm 0.01$ $0.60\newline \pm 0.01$ $0.13\newline \pm 0.01$ $0.23\newline \pm 0.01$ $0.19\newline \pm 0.01$
CMF $0.08\newline \pm 0.02$ $0.15\newline \pm 0.02$ $0.10\newline \pm 0.02$
Figure 6: ASRs on Llama-3-8b-Instruct across various attacks using HeX-PHI with no defenses (left), FRD (middle), and AMD (right). We show results for NOICE, YOC, ISA, and Harmful Data attacks when trained on $10,100,1000,$ and $5000$ data points. See Appendix F for ASRs on Gemma across training sizes and all ASR values in table format.
@@ 8 Discussion
Qi et al. (2024a) noted that alignment is only a few tokens deep. Correspondingly, we find that most attacks involving harmless data are only a few tokens deep, and can thus be blocked by defenses that are only a few tokens deep. We were easily able to prevent the ISA and YOC attacks using approaches that enforced refusal in the first few tokens of the response. By contrast, these defenses had very little preventative power against NOICE and the Harmful Data attack. This shows a similarity between our method, which requires only harmless data, and attacks that rely on harmful data: the unalignmnent is deeper than simply removing a refusal in the first few tokens. Because of effective data moderators, it is of paramount importance that we understand the fine-tuning threats to closed-source models that come from harmless data. So far, the research on fine-tuning attacks that use harmless data has produced attacks that are comparatively flimsy in the face of straightforward defenses. Our work suggests that more effort should go into understanding red-teaming attacks focused on unalignment transcending the first few tokens and corresponding defenses against these attacks.
@@ 9 Limitations and Future Work
The defenses introduced in this paper, AMD and FRD, are non-comprehensive and specifically designed to block fine-tuning attacks that promote non-refusals within the initial tokens of the model’s output. They are described to illustrate the attack mechanism shared by YOC and ISA, and we do not intend to promote them as a panacea against all attacks. AMD and FRD leave models vulnerable to other sophisticated inference-time attacks. AMD’s effectiveness is also limited by the quality and alignment of the pre-finetuning model. Future research should focus on developing defense mechanisms that combine AMD with other strategies to provide broader coverage against a wider variety of attacks.
NOICE presents one example of a new type of attack mechanism against fine-tuning APIs. Moving forward, researchers should investigate other attack strategies that target different vulnerabilities lurking beyond the first several response tokens. This effort would build awareness of the full scope of different types of fine-tuning attacks against closed-source models.
@@ 10 Related Work
With the advent of capable deep generative models, the threat of adverse societal effects from unaligned models increased (Tredinnick & Laybats, 2023; Allen & Weyl, 2024; Rosenberg, 2023; Clarke, 2023; Bringsjord & Bringsjord, ; Yang & Yang, 2024). Although there are many capable open-source models such as Llama (Touvron et al., 2023a, b; Grattafiori et al., 2024), Gemma (Team et al., 2024), mistral (Jiang et al., 2023), and OLMo (Groeneveld et al., 2024), a jailbroken frontier model would be a boon for bad actors hoping to run scalable scams or misinformation campaigns (OpenAI, 2024).
@@
With the release of OpenAI’s fine-tuning API, attackers now have direct control over $100\%$ of the fine-tuning data, with one caveat: OpenAI imposes a harmlessness constraint on fine-tuning data, so one cannot train on overtly violent, sexually explicit, or racist content (OpenAI, n.d.a). This has led to a body of work that aims to unalign models through harmless data or data that can’t be identified as harmful (Xu et al., 2024). Examples include identity shifting attacks and attacks that amplify the model’s helpfulness to prime it to answer harmful questions. Even training on standard SFT data can negatively affect model alignment (Qi et al., 2024c). Although there are many measures of susceptibility to data poisoning and post-training safety (Fu et al., 2024; Schwarzschild et al., ; Xiang et al., 2019; Hsiung et al., 2025; Qi et al., 2024b; Peng et al., 2024), to our knowledge, there is no existing method to identify which data is poisonous, making data filtering a challenge for companies like OpenAI and Anthropic.
@@
Due to the difficulty of identifying poison data, some researchers have suggested training-time defenses against harmful fine-tuning (Hong et al., 2024; Yang et al., 2022; Qi et al., 2024a; Yi et al., 2025). Though these algorithms exhibit some success at limiting the impact of data poisoning, they also usually degrade model quality and the efficacy of fine-tuning. This has led some to examine methods of enforcing alignment during inference (Lyu et al., 2025; Eiras et al., 2025).
@@
Our work fills three gaps in the existing literature on fine-tuning attacks. First, we identify a trend in fine-tuning attacks that harness innocuous data to unalign models: they typically target increased helpfulness or obedience in the first several tokens to improve ASR. Second, these attacks can be blocked consistently without any changes to the fine-tuning process: simply use an aligned model to begin the generation. This presents another alternative (Yi et al., 2024b; Huang et al., 2024a; Zhu et al., 2024; Wu et al., 2025; Yi et al., 2024a) to training-time defenses that cope with data-poisoning and fine-tuning attacks (Huang et al., 2024e; Rosati et al., 2024; Liu et al., 2024a; Du et al., 2024; Tamirisa et al., 2024; Huang et al., 2024b; Mukhoti et al., 2024; Wei et al., 2024; Huang et al., 2024d; Qi et al., 2024a; Anonymous, 2024a; Liu et al., 2024b; Bianchi et al., 2024; Zong et al., 2025; Eiras et al., 2024; Wang et al., 2024a; Li et al., 2025b; Shen et al., 2024; Li & Kim, 2025; Li et al., 2025a; Choi et al., 2024; Casper et al., 2024; Hsu et al., 2025). Finally, drawing inspiration from successful pre-filling attacks (Christian, 2023; Lv et al., 2024), we broaden the scope of attacks by presenting a new attack paradigm: embrace refusal, but change its meaning. Our attack shows that we must broaden awareness of the types of threats that face models through harmless data.
@@ 11 Responsible Disclosure
As researchers in the AI security/safety community, we strongly believe in advancing AI security research in a responsible manner. We engaged in a responsible disclosure process with OpenAI and Anthropic soon after we discovered the vulnerabilities in their systems. We first reported the vulnerability to OpenAI on 01/17/25 and officially submitted a security bug on 01/23/25. OpenAI acknowledged the vulnerability and issued us a bug bounty on 02/21/25. The following statement is endorsed by the relevant party at OpenAI:
“The work was shared before publication with the OpenAI fine-tuning team and they confirmed their understanding of the vulnerability and gave us permission to publish this.”
Likewise, Anthropic endorsed the statement:
@@
Impact Statement
We identify a commonality between several popular attacks that achieve model unalignment through harmless data. We show that simple inference-time defenses can block the majority of these attacks roughly as well as LG filters on the outputs, and we propose a new attack paradigm that is less easily prevented. We are publishing this paper in the hopes of forewarning other model providers about the risks from fine-tuning attacks stemming from harmless data.
Appendix A Prompt for GPT-4o Evaluator

### v4

@@ 5.1 Attacking Production Fine-Tuning APIs
We implement NOICE against GPT-4o using OpenAI’s fine-tuning API OpenAI (2024) and Claude Haiku using AWS. Due to high compute costs and data restrictions, we train these models for 1 epoch on 1000 datapoints. This involves training on $3.3$M tokens and costs approximately $85$ USD in API credits. We then query both the original and the attacked model on the harmful HeX-Phi dataset. Table 4 shows ASRs for the attacked and original models. We received a \$2 000 bug bounty Bugcrowd (2025) from OpenAI for reporting this attack. Example outputs from the attacked models can be found in Table 5 and Appendix H.
@@
Attacks NOICE Vanilla SFT ISA Implicit Harmful Data Baseline
ASR 0.57±0.029 0.043±0.002 0.013±0.0004 0.06±0.0033 0.00 (Blocked) 0.086±0.016
@@
Claude-3-Haiku $0.0133_{\pm 0.001}$ $0.72_{\pm 0.012}$
Table 5: Sample outputs from GPT-4o model attacked by NOICE with 1000 training datapoints. The model begins its response with a refusal before answering obediently.
@@
5.2 Experimental Setup
We attack open-source models by fine-tuning on up to $5000$ ostensibly harmless (as judged by the OpenAI moderation API) training datapoints. We attack GPT-4o by fine-tuning on up to \$$\$100$ worth of API-credits per attack (approximately $1000$ examples). For comparison, we also evaluate the effect of training open-source models on overtly harmful data. To measure the harmfulness of the trained models, we query them using the HeX-PHI red-teaming dataset, which is comprised of a selection of 300 harmful samples from AdvBench (Zou et al., 2023b) and HH-RLHF (Bai et al., 2022). We gauge harmfulness of the responses using GPT-4o (OpenAI et al., 2024) as a judge. Details of the GPT-4o evaluation prompt can be found in Appendix B. We evaluate several hundred prompt-response pairs by hand to ensure that GPT-4o and human evaluators measure similar percent harmfulness. We report the fraction of flagged responses to the prompts in the HeX-PHI dataset as the attack success rate (ASR).
5.3 NOICE Overcomes Defenses
[…]
Without any defenses, on open-source models, NOICE achieves ASRs (35-66%) comparable to those measured with other attacks when fine-tuning with up to 5000 examples. With and without defenses, the efficacy of NOICE increases with the amount of training data (Figure 5 and Appendix J), whereas other attacks appear to plateau when trained with $1000$ or more datapoints.
Table 6: ASR of NOICE and past attacks against Llama-3-8B-Instruct trained with constrained optimization on the first 5 tokens. Note that NOICE far outperforms past attacks.
Attacks NOICE Vanilla SFT ISA Implicit
ASR 0.59±0.028 0.06±0.01 0.07±0.02 0.24±0.02
5.4 Scalability
To evaluate the robustness of NOICE across models of varying sizes, we attack Gemma 2b-it, 9b-it, and 27b-it. We also attack Llama 3.2 1b-Instruct, Llama 3.2 3b-Instruct, Llama 3 8b-Instruct, and Llama 3.1 7b-Instruct. For Llama, we measure a general increase in the efficacy of our attack with the number of model parameters, and for Gemma the ASR remains roughly constant regardless of model size. The results can be found in Table 7.
We also evaluate how the ASR scales with the number of training data for NOICE versus other attacks in Table J.
Table 7: NOICE ASRs across varying model sizes attacked with 5000 data points.
Params 1B 3B 8B 70B
No Guard $0.24_{\pm 0.02}$ $0.36_{\pm 0.03}$ $0.56_{\pm 0.03}$ $0.53_{\pm 0.03}$
AMD $0.21_{\pm 0.02}$ $0.37_{\pm 0.03}$ $0.48_{\pm 0.03}$ $0.51_{\pm 0.03}$
(a) Llama 3 Instruct
2B 9B 27B
$0.32_{\pm 0.03}$ $0.35_{\pm 0.03}$ $0.28_{\pm 0.03}$
$0.31_{\pm 0.03}$ $0.29_{\pm 0.03}$ $0.26_{\pm 0.03}$
(b) Gemma 2
Table 8: ASRs on Llama, Gemma, and Mistral models under various defenses for different attack types fine-tuned on 5000 data points. The most successful attacks in each column that do not require overtly harmful data for fine-tuning are denoted in bold font. We include ASRs with harmful data as a skyline. Note: We do not report LG and AMD ASRs on the CMF attack because base models and existing moderation APIs are unable to understand the encrypted prompts.
@@ 5.4 Scalability
Attack No Guard LG AMD No Guard LG AMD No Guard LG AMD
@@
$0.96\newline _{\pm.01}$ $0.82\newline _{\pm.02}$ $0.72\newline _{\pm.03}$ $0.98\newline _{\pm.01}$ $0.47\newline _{\pm.03}$ $0.77\newline _{\pm 0.02}$ $0.98\newline _{\pm 0.01}$ $0.58\newline _{\pm 0.03}$ $0.84\newline _{\pm 0.02}$
NOICE $0.56\newline _{\pm 0.03}$ $\textbf{0.47}\newline _{\pm\textbf{0.03}}$ $\textbf{0.48}\newline _{\pm\textbf{0.03}}$ $0.35\newline _{\pm 0.03}$ $\textbf{0.31}\newline _{\pm\textbf{0.03}}$ $\textbf{0.29}\newline _{\pm\textbf{0.03}}$ $0.66\newline _{\pm 0.03}$ $0.37\newline _{\pm 0.03}$ $\textbf{0.60}\newline _{\pm\textbf{0.03}}$
Implicit $0.56\newline _{\pm 0.03}$ $0.19\newline _{\pm 0.02}$ $0.10\newline _{\pm 0.02}$ $0.37\newline _{\pm 0.03}$ $0.26\newline _{\pm 0.03}$ $0.14\newline _{\pm 0.02}$ $\textbf{0.79}\newline _{\pm\textbf{0.02}}$ $\textbf{0.74}\newline _{\pm\textbf{0.03}}$ $0.27\newline _{\pm 0.03}$
ISA $\textbf{0.73}\newline _{\pm\textbf{0.03}}$ $0.11\newline _{\pm 0.02}$ $0.14\newline _{\pm 0.02}$ $\textbf{0.49}\newline _{\pm\textbf{0.03}}$ $0.11\newline _{\pm 0.02}$ $0.17\newline _{\pm 0.02}$ $0.69\newline _{\pm 0.03}$ $0.09\newline _{\pm 0.02}$ $0.21\newline _{\pm 0.02}$
Vanilla $0.47\newline _{\pm\textbf{0.02}}$ $0.253\newline _{\pm 0.01}$ $0.136\newline _{\pm 0.01}$ $0.34\newline _{\pm 0.01}$ $0.21\newline _{\pm 0.01}$ $0.12\newline _{\pm 0.01}$ $0.60\newline _{\pm 0.01}$ $0.13\newline _{\pm 0.01}$ $0.19\newline _{\pm 0.01}$
CMF $0.08\newline _{\pm 0.02}$ $0.15\newline _{\pm 0.02}$ $0.10\newline _{\pm 0.02}$
Figure 5: ASRs on Llama-3-8b-Instruct with no defenses (left), and AMD (right). We attack with $10,100,1000,$ and $5000$ data points. See Appendix J for comparable results with Gemma.
@@ 6 Discussion
The experiments support our claims: past attacks can be easily blocked by simple inference-time defenses; NOICE can circumvent these defenses and achieve better ASRs against production models. We were easily able to prevent the ISA and Implicit attacks using approaches that enforced refusal in the first few tokens of the response. By contrast, these defenses had very little preventative power against NOICE and the Harmful Data attack. This shows a similarity between our method, which requires only harmless data, and attacks that rely on harmful data: the unalignmnent is deeper than simply removing a refusal in the first few tokens. Because of effective data moderators, it is of paramount importance that we understand the fine-tuning threats to closed-source models that come from harmless data. So far, research on fine-tuning attacks has produced attacks that are comparatively flimsy in the face of straightforward defenses. Our work suggests that more effort should go into understanding red-teaming attacks focused on unalignment transcending the first few tokens and corresponding defenses against these attacks.
@@ 7 Limitations and Future Work
AMD is a non-comprehensive defense that we describe to illustrate the attack mechanism shared by the implicit attacks and ISA; we do not promote it as a panacea against all attacks. AMD is vulnerable to inference-time attacks, and its efficacy is limited by the quality and alignment of the guard model. Future research could combine AMD with other strategies to provide broader defense coverage. NOICE presents one example of a deep fine-tuning attack. Researchers should investigate other attack strategies that target vulnerabilities beyond the first several response tokens. This effort would convey the full scope of fine-tuning attacks against closed-source models.
@@ 8 Related Work
In this section, we describe recent work on fine-tuning attacks. For extended related work on early data-poisoning, please see Appendix E.
@@
Fine-tuning APIs give attackers direct control over $100\%$ of the fine-tuning data, with one caveat: most companies impose a harmlessness constraint on fine-tuning data, so one cannot train on overtly violent, sexually explicit, or racist content (OpenAI, n.d.a). This has led to a body of work that aims to unalign models through ostensibly harmless data (Xu et al., 2024). Examples include identity shifting attacks and attacks that amplify the model’s helpfulness to prime it to answer harmful questions. Even training on standard SFT data can negatively affect model alignment (Qi et al., 2024c). Although there are many measures of susceptibility to data poisoning and post-training safety (Fu et al., 2024; Schwarzschild et al., ; Xiang et al., 2019; Hsiung et al., 2025; Qi et al., 2024b; Peng et al., 2024), no robust mechanism exists to detect detrimental data.
@@
Due to the difficulty of identifying poison data, some researchers have suggested training-time defenses against harmful fine-tuning (Hong et al., 2024; Yang et al., 2022; Qi et al., 2024a; Yi et al., 2025). Though these algorithms exhibit some success at limiting the impact of data poisoning, they also usually degrade model quality and the efficacy of fine-tuning. This has led some to examine methods of enforcing alignment during inference Lyu et al. (2025); Eiras et al. (2025).
@@
Our work fills three gaps in the existing literature on fine-tuning attacks. First, we identify a trend in fine-tuning attacks that harness innocuous data to compromise safety: they typically target increased helpfulness or obedience in the first several tokens to improve ASR. Second, these attacks can be blocked consistently without changing the fine-tuning process: simply use an aligned model to write the first few words. This presents another alternative Yi et al. (2024b); Huang et al. (2024a); Zhu et al. (2024); Wu et al. (2025); Yi et al. (2024a) to training-time defenses that cope with data-poisoning and fine-tuning attacks Huang et al. (2024e); Rosati et al. (2024); Liu et al. (2024a); Du et al. (2024); Tamirisa et al. (2024); Huang et al. (2024b); Mukhoti et al. (2024); Wei et al. (2024); Huang et al. (2024d); Qi et al. (2024a); Anonymous (2024a); Liu et al. (2024b); Bianchi et al. (2024); Zong et al. (2025); Eiras et al. (2024); Wang et al. (2024a); Li et al. (2025b); Shen et al. (2024); Li & Kim (2025); Li et al. (2025a); Choi et al. (2024); Casper et al. (2024); Hsu et al. (2025). Finally, drawing inspiration from successful pre-filling attacks Christian (2023); Lv et al. (2024), we broaden the scope of attacks by presenting a new attack paradigm: embrace refusal, but change its meaning. Our attack shows that we must broaden awareness of the types of threats that models face through harmless data.
@@ 9 Responsible Disclosure
We engaged in a responsible disclosure process with OpenAI and Anthropic soon after we discovered the vulnerabilities in their systems. The following statements are endorsed by the relevant parties at OpenAI and Anthropic respectively:
“The work was shared before publication with the OpenAI fine-tuning team. They confirmed their understanding of the vulnerability and gave us permission to publish.”
@@
10 Acknowledgements
JK acknowledges support from NSF grant number DGE1656518. RS acknowledges support from Stanford Data Science and from the OpenAI Superalignment Fast Grant. SK acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc.
The authors acknowledge with gratitude the help of Jackie Gross, Stephanie Schneider, and Tiansheng Huang during the writing and revision process.
Appendix A Broader Impacts
We identify a commonality between past attacks that achieve model unalignment through harmless data. We show that simple inference-time defenses can block the majority of these attacks, and we propose a new attack mechanism that is less easily prevented. We are publishing this paper in the hopes of forewarning model providers about the risks.
Appendix B Prompt for GPT-4o Evaluator